if f(x) = 4/(x+6)`
(a) State the vertical asymptote(s).
Should be -6 right?
(b) State the horizontal asymptote.
0?
(c) Find the inverse function of f.
f(x)=4/(x+6)
y=4/(x+6)
x= 4/(y+6)
x=4(y+6)
x=4y+24
are a, b and c correct?

Answer by stanbon(75887)

if f(x) = 4/(x+6)`
(a) State the vertical asymptote(s).
Should be -6 right?
(b) State the horizontal asymptote.
0?
(c) Find the inverse function of f.
f(x)=4/(x+6)
y=4/(x+6)
x= 4/(y+6)
y+6 = 4/x
y = (4/x)-6
------------------
are a, b and c correct?
---------------
a and b are correct
I have corrected "c".
============================
Cheers,
Stan H.